OpenStudy (anonymous):

NEED HELP PLEASE!!!!!!!! write the equation of the line that is parallel to the line y = -3x + 12, and passes through the point (-1, 6)

OpenStudy (majesty69):

y=-3x+9

OpenStudy (anonymous):

the answers are A. y = 1/3x + 7 B. y = -3x + 3 C. y = 1/3x + 3 D. y = -3x + 7

OpenStudy (anonymous):

@Majesty69 i forgot to put the answer choices. I always do that.

OpenStudy (majesty69):

hmmmmmmmm sorry i made a small mistake :( it is B.

OpenStudy (anonymous):

thanks. every time i post a question on here i always forget to put the answer choices :/

OpenStudy (majesty69):

hahaha its okay :)